OpenStudy (anonymous):

review the steps to solve the problem. What is wrong in these steps? 4* (x+2) = 14 Step 1: 4x + 2 = 14 Step 2: 4x = 12 Step 3: x=3 I think that the first error is in step one. Shouldn't it be 4x + 8 = 14... or is this all correct

OpenStudy (anonymous):

Yes it should be 4x+8=14

OpenStudy (anonymous):

so then it would be 4x+8=14 4x=6 then what would x be?

OpenStudy (anonymous):

x would then be \[\frac{ 6 }{ 4 } or \frac{ 3 }{ 2 }\] when simplified.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No Problem, Good Luck :)